OpenStudy (anonymous):

How do you do this problem, i need a refresher haha. 3x-2-x+6=24

OpenStudy (anonymous):

3x-2-x+6=24 3x-x-2+6=24 2x+4=24 2x=24-4 2x=20 x-20/2=10

OpenStudy (dumbcow):

what they said ^^ combine like terms isolate the variable by dividing the coefficient